SOLVE AND SHOW COMPLETE SOLUTION



1. if y= ax^3 + bx^2 and its point of inflection is at (2,8), what are the values of a and b?




2. graph the curve and find the point of inflection: y= (9x^2 - x^3 + 6)/ (6).




3. graph the curve and find the point of inflection: y= x^3 - 3x^2 + 6.

1
Expert's answer
2021-12-27T16:17:31-0500

1.


"y= ax^3 + bx^2"

"y'=3 ax^2 +2 bx"

"y''=6ax+2b"

"y''(2)=0=>6a(2)+2b=0=>b=-6a"

"y(2)=8=>8a+4b=8=>b=-2a+2"

"-6a=-2a+2=>a=-\dfrac{1}{2}"

"b=-6(-\dfrac{1}{2})=3"

"y=-\dfrac{1}{2}x^3+3x^2"

"a=-\dfrac{1}{2}, b=3"

2.


"y= \dfrac{9x^2 - x^3 + 6}{6}"

"D(y): (-\infin, \infin)"


"E(y): (-\infin, \infin)"


"y\to\infin" as "x\to -\infin"

"y\to-\infin" as "x\to \infin"


"y(0)=1"



"y'= \dfrac{18x - 3x^2 }{6}= 3x-\dfrac{x^2}{2}"

Find the critical number(s)


"y'=0=>3x-\dfrac{x^2}{2}=0"

"x_1=0, x_2=6"

Critical numbers: "0, 6."


"y(0)=1"

"y(6)= \dfrac{9(6)^2 - (6)^3 + 6}{6}=19"

If "x<0, y'<0, y" decreases.

If "0<x<6, y'>0, y" increases.

If "x>6, y'<0, y" decreases.

The function "y" has the local maximum with value of "19" at "x=6."

The function "y" has the local minimum with value of "1" at "x=0."



"y''=3-x"

"y''=0=>3-x=0=>x=3"

"y(3)= \dfrac{9(3)^2 - (3)^3 + 6}{6}=10"

If "x<3, y''>0, y" is concave up.

If "x>3, y''<0, y" is concave down.

Point "(3, 10)" is the inflection point.

Graph the function




3.


"y= x^3 - 3x^2 + 6"

"D(y): (-\infin, \infin)"


"E(y): (-\infin, \infin)"


"y\to-\infin" ​as"x\to -\infin"

"y\to\infin" ​as"x\to \infin"


"y(0)=6"



"y'=3x^2-6x"

Find the critical number(s)


"y'=0=>3x^2-6x=0"

"x_1=0, x_2=2"

Critical numbers: "0, 2."


"y(0)=6"

"y(2)=(2)^3 - 3(2)^2 + 6=2"

If "x<0, y'>0, y" increases.

If "0<x<2, y'<0, y" decreases.

If "x>2, y'>0, y" increases.

The function "y" has the local maximum with value of "6" at "x=0."

The function "y" has the local minimum with value of "2" at "x=2."



"y''=6x-6"

"y''=0=>6x-6=0=>x=1"

"y(1)=(1)^3 - 3(1)^2 + 6=4"

If "x<1, y''<0, y" is concave down.

If "x>1, y''>0, y" is concave up.

Point "(1, 4)" is the inflection point.

Graph the function